Prev Next

SW&IT Meeting Minutes 2018-12-10

Attendees: David, Igor, Matthieu, Petr, Stephen, Yang

Outlook

Labor Reporting

Report auto-generated every, Do not edit this table, use entries in smartsheets.

20181119
LR20181119 - Bec, Matthieu 64.0
LR20181119 - Colson, Andrew 78.0
LR20181119 - Edmeades, David 64.0
LR20181119 - Sola, Igor 53.0
LR20181119 - Kubanek, Petr 64.0
LR20181119 - Hooper, Stephen 64.0
LR20181119 - Zhang, Yang 40.0
LR20181119 - Zhang, Xianyu 0.0

20181203
LR20181203 - Bec, Matthieu 41.0
LR20181203 - Colson, Andrew 32.0
LR20181203 - Edmeades, David 8.0
LR20181203 - Sola, Igor 24.0
LR20181203 - Kubanek, Petr 40.0
LR20181203 - Hooper, Stephen 80.0
LR20181203 - Zhang, Yang 40.0
LR20181203 - Zhang, Xianyu 0.0

Transition plan

Handover steps

  1. primary + secondary do self learning, gather information, read wiki
  2. primary organize teamviewer/zoom walk through with Kellee and secondary
  3. primary draft techtalk w. input from secondary
  4. techtalk review by Kellee
  5. primary does a team presentation

Please update this table to indicate progress.


Application(s) Primary Secondary WeekSorted ascending
ALH Igor(2)
Matthieu (2)
1
AzCam
AC (2)
Petr (2)
2
Weatherstation
AC (2)
Petr (2)
2
Status Server Stephen (2)
Matthieu (2)
2
MODS
Yang (3)
Igor (1/2) 2
INDI Igor (2)
Matthieu (2)
3
Telemetry cleanup Stephen (1) Petr 3
Telem-VIZ
Yang (1)
Stephen (1) 3
FACSUM,WEATHERSTATION
AC (2)
Matthieu (2)
4
particle monitors AC Petr 4

News from management

  • PEPSI/pol high profile run in January: AGW3 & 4 must be ready
  • ARGOS half-night Dec 22 - confirmed

Previous Action Items

  • Skill set list (DE)

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

  • network dropout - antenna monitoring?
  • iLocater Single-Mode Data Fibers - CRA/CRB to Electronics Shop
  • BCU f/w upgrade documentation
  • revert simbad harvard-strasbourg /et/hosts hack from Friday
  • Dust monitor status
  • AO improvement: Luci kwds
  • AO cabinet switch upgrade - follow up
  • wifi summit upgrade (note: end-of-year deadline)
  • Tucson CR power overhaul plan (note: end-of-year deadline)
  • n507/obs upgrade (note: end-of-year deadline)
  • synology new drives
  • PGray mirror cover req. feedback
  • Deployment on page https://wiki.lbto.org/Software/ArgosWiki broken
  • Argos Dec 22, second half - TBC
  • Installing python packages on flao-sxwfs

This week in github

Team reports

YZ

  • IT issue #7285 (MODS1 data machine freeze problem), checked dmesg and several logs with no obvious problems, forwarding it to Rick Pogge for further investigation.
  • IT issue #7500 (binocular LBCR + MODS1 observing failed with GetRotatorTrajector error), Petr suggested PCS problem and will investigate.
  • IT issue #6395 (MODS naming/print statements as seen in UI). Olga confirmed it was an OT issue not related to MODS.
  • MODS data computer high storage warning resolved (Rick Pogge removed some old data under /lhome/data).
  • Obtained passwords for MODS computers, need to update the shared password document next.
  • Working on AO Cravat web interface code:
    • Python code convert from fits to png and then thumbnail. Currently used a combination of astropy tools and own code. Maybe better to produce a single Python code.
    • Working on Polymer/JS9 integration, some progress, still needs debugging and tidying things up for a demo.
  • Working on MODS wiki documents.

IS

This Week
  • LUCI
    • Working on AFC debug tool. The tool copies AFC subwindow files as they get generated and display them locallly in a python program. The program applies the same logic LUCI does in the backend.
  • OT
    • Debugging #7499. It was reported as a scripting issue but it turned out to be user error.
Next week
  • LUCI
    • Finish AFC debug tool
  • OT
    • Possibly ARGOS/ESM binocular mode

PK

  • PCS copointing fix
  • work underway improving AGW OAC code for the new housekeeping units
  • LBC NS guiding meeting
next:

AC

  • Fixed ARGOS wiki
  • ARGOS handover items: spare computer, open and new ITs
  • OAC AGW RPC error
  • oaccontrol code refactoring (see PK)
  • AGW test procedure

MB

  • OC, WPC, management
  • ovms-tt weekly, ordered IPC x3
  • IMS fix layout bugs
  • review PG m1 cover req document (on-going)
  • LUCI/AO keyword recap spreadsheet
  • OAC discussion
next:
  • SMT/wifi meeting
  • flao-sx python3 Fabio
  • VT meet up?
  • iLocater / 3L / spectro PDR

DE

  • Autodesk software discussion
    • Decided that IT group rather than Ketiv should install client software for our users
    • Standardizing on 2018 versions of main products; other versions do not play well with 2018 Vault server
    • Started updating at users' request
  • Conference room audio
    • Preliminary results positive
  • Laura hardware issues ongoing, replacement hardware did not solve the issue for long.
  • AGw feature freeze? I'd like to image the current drive for easy replication
  • Rep change at CDW; asked new person for update on 12TB drives this morning.
  • Will send CV
  • Thoughts on power plan?

SH

  • continued working on Wifi summary
  • history information to http://info (not great yet, but does mostly work).
  • continuing mountain network woes.
  • iLocater network and cabinets
  • Ordered computers/cards for obs replacement.
  • Ordered laptop for Bucki
next:
  • AO cabinet switch configuration (try to feed 100Mbps in the lab).
  • Microwave information collection
  • iLocater
  • refine info page -- add info, better history, less bandwidth, mobile site refinement -- not very important so lower priority.
  • DIMM computer replacement -- quotes in, but discuss with involved parties.
  • ARGOS computer spare.
  • continued mountain link fun I am sure.
  • LBTI computer IP addresses -- check back with Paul (on NTP time as well).
Topic revision: r21 - 10 Dec 2018, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback